# See2Refine: Vision-Language Feedback Improves LLM-Based eHMI Action Designers

### v1

@@ Abstract
Automated vehicles lack natural communication channels with other road users, making external Human–Machine Interfaces (eHMIs) essential for conveying intent and maintaining trust in shared environments. However, most eHMI studies rely on developer-crafted message–action pairs, which are difficult to adapt to diverse and dynamic traffic contexts. A promising alternative is to use Large Language Models (LLMs) as action designers that generate context-conditioned eHMI actions, yet such designers lack perceptual verification and typically depend on fixed prompts or costly human-annotated feedback for improvement. We present See2Refine, a human-free, closed-loop framework that uses vision–language model (VLM) perceptual evaluation as automated visual feedback to improve an LLM-based eHMI action designer. Given a driving context and a candidate eHMI action, the VLM evaluates the perceived appropriateness of the action, and this feedback is used to iteratively revise the designer’s outputs, enabling systematic refinement without human supervision. We evaluate our framework across three eHMI modalities (lightbar, eyes, and arm) and multiple LLM model sizes. Across settings, our framework consistently outperforms prompt-only LLM designers and manually specified baselines in both VLM-based metrics and human-subject evaluations. Results further indicate that the improvements generalize across modalities and that VLM evaluations are well aligned with human preferences, supporting the robustness and effectiveness of See2Refine for scalable action design.  [^1] The source code, prompts, Blender scenarios, and rendered clips are available at https://anonymous.4open.science/r/eHMI_LLM_finetuning-3A74
@@
Ding Xia1, Xinyue Gui1, Mark Colley1,2, Fan Gao1, Zhongyi Zhou1,3 Dongyuan Li1, Renhe Jiang1, Takeo Igarashi1 1The University of Tokyo, 2University College London, 3Google {dingxia1995, fangao0802, lidy94805}@gmail.com, zhongyizhou@google.com gui-xinyue@g.ecc.u-tokyo.ac.jp, m.colley@ucl.ac.uk, takeo@acm.org
@@ 1 Introduction
Automated vehicles (AVs) are expected to reach Level 4 autonomy within a decade Agrawal et al. (2023); Chen et al. (2023). Without human drivers, AVs usually lack effective communication channels with other road users, such as pedestrians, cyclists, and human drivers Fagnant and Kockelman (2015). This gap causes confusion and distrust, leading to traffic issues Colley et al. (2025). External Human-Machine Interfaces (eHMIs) are emerging as promising solutions to address this communication challenge Dey et al. (2020a). eHMIs can comprise devices such as mechanical eyes Chang et al. (2022), mechanical arms Gui et al. (2024b), and lightbars Dey et al. (2020b) mounted on the exterior of AVs. These devices serve as communication interfaces that convey messages like “I am stopping”. For instance, an eHMI can effectively express directional intent through eye gazing, finger pointing, or illuminating particular sections of a lightbar Gui et al. (2024a).
With eHMIs, researchers typically design eHMI interactions via predefined actions, limiting adaptability in dynamic and unpredictable environments Dey et al. (2020a); de Winter and Dodou (2022). To address this rigidity, LLMs offer a data-driven alternative by generating context-dependent eHMI actions from traffic scenarios, reducing the reliance on manually defined rules Radford et al. (2019). Moreover, prior work shows that, when equipped with carefully crafted system prompts, LLM-based action designers can perform at a level comparable to that of human designers Xia et al. (2025). However, text-based prompts can only capture a portion of the details involved in eHMI installation and are unable to convey rich visual information, such as sizes, speeds, and observation angles, which are essential for designing actions in dynamic scenarios Tellex et al. (2020); Cao et al. (2024); Majumder et al. (2024).
Using human feedback to guide the extraction of knowledge related to specific types of eHMI serves as a viable solution. Reinforcement Learning from Human Feedback (RLHF) Bai et al. (2022a) has proven effective in developing LLMs at the expert-level in specific tasks. However, applying RLHF to the eHMI LLM designer involves extensive human annotations, similar to other RLHF studies Kaufmann et al. (2023); Jin et al. (2023); Wang et al. (2025). Reinforcement Learning from AI Feedback (RLAIF) offers an alternative approach by leveraging visual perception capabilities of Vision-Language Models (VLMs) to provide judgments that are close to human-level Lee et al. (2024a); Lu et al. (2024). To develop a framework capable of designing, evaluating, and self-improving in a cost-effective, rapid, and automated way, we pose the following research question:
@@
In this work, we propose See2Refine that features LLM designers to autonomously design, evaluate, and iteratively refine eHMI actions, utilizing visual perception from VLMs as automated feedback in a cost-effective manner. In our experiments, we trained separate DesignerLLM models for each of the three eHMI modalities: lightbar, eyes, and arm. After three rounds of iterative learning, the DesignerLLM models exhibited a clear alignment in preferences with VLM raters. Additionally, our sampling method to expand the action database effectively reduces annotation costs and time without compromising performance. Subsequently, we recruited 18 participants to evaluate the eHMI actions generated by five LLMs: two 7B models (one base model, one DesignerLLM) and three state-of-the-art models. Extensive results show that utilizing the visual perception of VLMs can also enhance human evaluation scores. The main contributions can be summarized as follows:
  • We propose a new See2Refine framework that integrates VLM-based perceptual evaluation as automated visual feedback to improve LLM-based eHMI action designers (DesignerLLM).
  • We discuss methods for effectively expanding the action database while maintaining equivalent VLM preference alignment performance.
  • We conduct a user study to demonstrate that aligning with VLM raters’ preferences can also enhance human evaluation scores.
2 Related Works
@@
Existing eHMI action-planning mostly follows predefined, designer-authored rules per modality: text/icons use traffic-rules messages Eisele and Petzoldt (2022); Eisma et al. (2021); color/light use flashing patterns based on intuitive color associations Bazilinskyy et al. (2019); Dey et al. (2020b); and anthropomorphic cues (e.g., eyes, arm gestures) draw on nonverbal communication principles Mahadevan et al. (2018); Ochiai and Toyoshima (2011). While effective, these manual designs are labor-intensive and hard to scale to complex real-world scenarios Gui et al. (2023); de Winter and Dodou (2022). Recent work uses LLMs for eHMI action planning via prompt engineering Xia et al. (2025), but this remains prompt-dependent and largely task-level, lacking mechanism-level understanding, requiring more advanced methods for higher-quality action planning.
@@
Reinforcement Learning from AI Feedback (RLAIF) replaces the collection of human preferences in RLHF by using stronger teacher models to generate reward signals Bai et al. (2022b); Lee et al. (2023). In multimodal settings, VLMs can be fine-tuned as rubric-following judges for images/videos with human-level agreement competitive with GPT-4V Lee et al. (2024a); Lu et al. (2024); Chen et al. (2024). Systems such as Prometheus-Vision, WildVision-Arena/Bench, and MLLM-as-a-Judge show that VLM judges can follow instructions and output pointwise/pairwise scores that preserve human preference orderings Lee et al. (2024a); Lu et al. (2024); Chen et al. (2024), while VHELM and LLaVA-Critic further standardize multimodal evaluation and reduce reliance on closed models Lee et al. (2024b); Xiong et al. (2025). Multimodal variants of RLAIF/RLHF-V use AI-generated preferences to train reward models and optimize LVLM policies with minimal human labeling Ahn et al. (2024); Yu et al. (2025); Sun et al. (2023). For eHMI, VLM judges correlate well with human ratings on LLM-generated action clips, enabling closed-loop optimization where an LLM proposes actions, a VLM scores them, and updates (e.g., policy gradient or DPO-style) refine the space without human studies Wang et al. (2024).
@@ 3 Method
As shown in Figure 2, our See2Refine framework contains (1) Scenario-Message Pair Generation, (2) Model Asset & Action Rendering, (3) Multi-Metric Evaluation System, (4) Format-aware Fine-tuning, and (5) Iterative Preference-based Learning.
@@ 3.1 Scenario-Message Pair Generation
We first generate traffic scenarios by combining six different factors, inspired by existing eHMI scenario studies Dey et al. (2020a); Colley and Rukzio (2020). The factors we consider are listed below. By combining these factors, we generate a total of 6,912 unique condition combinations.
  • Communication Relationships (4): 1st-/3rd-person perspective $\times$ one-to-one/one-to-many.
@@
  • Message Type (8): Instruction, advisory, question, answer, current, historical, predictive, warn.
@@
  • Safety Implications (3): Critical (0–5 m), moderate (5–10 m), routine (>>10 m).
Second, conditioned on these scenarios, we generate 20,736 intended messages containing scenario information using larger LLMs, including OpenAI GPT-4.1 OpenAI (2025), Claude 3.7 Sonnet Anthropic (2025), and Qwen3-235B-A22B Qwen Team, Alibaba Cloud (2025). To eliminate redundancy, we employ a sentence transformer Reimers and Gurevych (2019) to compare messages. We sample 70% of these messages using the farthest point sampling algorithm Eldar et al. (1997) to ensure sufficient variance. In addition, we conduct a small user-rating study on 200 randomly selected scenario–message pairs. Two eHMI researchers independently rate the validity of each pair on a 7-point Likert scale (1=Strongly Disagree, 7=Strongly Agree). The mean rating is 5.3 (SD=1.9), between “Somewhat Agree” (5) and “Agree” (6), suggesting that the generated scenarios were generally of high quality.
@@ 3.2.1 Model Development and Definition
In this study, we focus on three eHMI modalities: lightbar, eyes, and arm. The first two modalities represent classic eHMI designs Dey et al. (2020a); Deb et al. (2018); Rouchitsas and Alm (2019); Benderius et al. (2017); Colley et al. (2020) with simpler control mechanisms, while the arm modality requires managing more parameters, which presents additional challenges for DesignerLLM. In Blender (Blender Foundation, 2025), we implemented each modality in two message emitter types (self-driving cars and delivery robots), resulting in six 3D models (Figure 3 (a)). Modality descriptions and permitted actions are as follows:
@@
  • Light Bar is mounted on the vehicle front, slightly extending to both wings. It has 16 binary lighting regions (0=off, 1=on); e.g., “0000000011111111” denotes left-half off and right-half on (vehicle perspective). Brightness is fixed and color is cyan Dey et al. (2020b).
@@
We also specify transition speeds between action states. The eye and arm transitions range from 0.5 to 2.0s (0.1s steps), while the lightbar transitions range from 0.1 to 1.0s (0.1s per step). This enables smooth sequences while preserving temporal detail at the video FPS.
@@ 3.2.2 Rendering Pipeline
For rendering, we typically consider three conditions: camera direction, camera distance (relative to the object), and background scenes, as illustrated in Figure 3 (b). Regarding camera direction, aligned with Section 3.1, we create 12 cameras arranged in clockwise order. Regarding camera distance, which reflects the safety implications (Section 3.1), we position the cameras at three different distances. These positions can be accessed via Python to generate videos that simulate the perspectives of other road users. We use a plain background, unlike the previous work Xia et al. (2025), to avoid contextual cues and reduce rendering time. Removing scene backgrounds decreases rendering time to approximately $1/10$, while lower GPU memory usage also enables more parallel threads, further reducing wall-clock time by around 4× for batch rendering.
We render action videos using Blender version 4.5 Blender Foundation (2025) on a GPU-equipped device featuring four NVIDIA GTX Titan cards with 24 GB of RAM each, which allows for up to 16 concurrent threads. The average rendering time for 1,000 clips is approximately one hour, with each thread consuming around 1.1 GB of GPU RAM. Each frame is rendered at a resolution of 512×512 pixels, centered on the eHMI-equipped AV, as shown in Figure 3 (c). These clips are rendered at four FPS, with a maximum action change rate for each eHMI set to ensure that the frame rate captures all relevant details, following Nyquist-Shannon sampling theorem Nyquist (1928).
@@ 3.3 Multi-Metric Evaluation System
Recent studies show that VLMs demonstrate capabilities and preferences similar to those of humans Lee et al. (2024a); Lu et al. (2024); Chen et al. (2024). In our task, we aim to borrow knowledge from existing Human-Computer Interaction research Colley et al. (2020); Colley and Rukzio (2020); Colley et al. (2025) and use the features of VLMs to develop a multi-metric evaluation system.
@@ 3.4 Format-aware Fine-tuning
This phase trains DesignerLLM-Apprentice to reliably output eHMI-specific action formats, as illustrated in Figure 2. We fine-tune two base model sizes, Qwen2.5-7B-Instruct and Qwen2.5-1.5B-Instruct Qwen Team (2024) Qwen Team (2024) as “apprentices” for later optimization. Training data is sampled from the shared action database by selecting the candidate action with the best kernel score per scenario.

### v2

@@ Abstract
Automated vehicles lack natural communication channels with other road users, making external Human-Machine Interfaces (eHMIs) essential to convey intent and maintain trust in shared environments. However, most eHMI studies rely on developer-crafted message-action pairs, which are difficult to adapt to diverse and dynamic traffic contexts. A promising alternative is to use Large Language Models (LLMs) as action designers that generate context-conditioned eHMI actions, yet such designers lack perceptual verification and typically depend on fixed prompts or costly human-annotated feedback for improvement. We present See2Refine, a human-free, closed-loop framework that uses vision-language models (VLMs) for perceptual evaluation as automated visual feedback to improve an LLM-based eHMI action designer. Given a driving context and a candidate eHMI action, the VLM evaluates the perceived appropriateness of the action, and this feedback is used to iteratively revise the designer’s output, enabling systematic refinement without human supervision. We evaluate our framework across three eHMI modalities (lightbar, eyes, and arm) and multiple LLM model sizes. Across settings, our framework consistently outperforms prompt-only LLM designers and manually specified baselines in both VLM-based metrics and human-subject evaluations. The results further indicate that the improvements are generalized across modalities and that VLM evaluations are reasonably aligned with human preferences in our controlled settings, supporting the robustness and effectiveness of See2Refine for scalable action design.  [^1] The source code, prompts, and Blender scenarios are available at https://github.com/ApisXia/see2refine
@@
Ding Xia1, Xinyue Gui1, Mark Colley1,2, Zhongyi Zhou3, Fan Gao1 Dongyuan Li1thanks: Corresponding author., Renhe Jiang1, Takeo Igarashi1 1The University of Tokyo,  2University College London,  3Google dingxia1995@gmail.com, gui-xinyue@g.ecc.u-tokyo.ac.jp, m.colley@ucl.ac.uk, zhongyizhou@google.com, fangao0802@gmail.com {lidy, jiangrh}@csis.u-tokyo.ac.jp, takeo@acm.org
@@ 1 Introduction
Automated vehicles (AVs) are expected to reach Level 4 autonomy within a decade Agrawal et al. (2023); Chen et al. (2023). Without human drivers, AVs usually lack effective communication channels with other road users, such as pedestrians, cyclists, and human drivers Fagnant and Kockelman (2015). This gap causes confusion and distrust, leading to series traffic issues Colley et al. (2025). External Human-Machine Interfaces (eHMIs) are emerging as promising solutions to address this communication challenge Dey et al. (2020a). eHMIs can comprise devices such as mechanical eyes Chang et al. (2022), mechanical arms Gui et al. (2024b), and lightbars Dey et al. (2020b) mounted on the exterior of AVs. These devices serve as communication interfaces that convey messages like “I am stopping”. For instance, an eHMI can effectively express directional intent through eye gazing, finger pointing, or illuminating particular sections of a lightbar Gui et al. (2024a).
With eHMIs, researchers typically design eHMI interactions via predefined actions, limiting adaptability in dynamic and unpredictable environments Dey et al. (2020a); de Winter and Dodou (2022). To address this rigidity, LLMs offer a data-driven alternative by generating context-dependent eHMI actions from traffic scenarios, reducing the reliance on manually defined rules Radford et al. (2019). Moreover, previous work shows that, when equipped with carefully crafted system prompts, LLM-based action designers can perform at a level comparable to that of human designers Xia et al. (2025). However, text-based prompts can only capture a portion of the details involved in eHMI installation and are unable to convey rich visual information, such as sizes, speeds, and observation angles, which are essential for designing actions in dynamic scenarios Tellex et al. (2020); Cao et al. (2024); Majumder et al. (2024).
Using human feedback to guide the extraction of knowledge related to specific types of eHMI serves as a viable solution. Reinforcement Learning from Human Feedback (RLHF) Bai et al. (2022a) has proven effective in developing LLMs at the expert-level in specific tasks. However, applying RLHF to LLM-based eHMI action designers requires extensive human annotations, similar to other RLHF studies Kaufmann et al. (2023); Jin et al. (2023); Wang et al. (2025). Reinforcement Learning from AI Feedback (RLAIF) offers an alternative approach by leveraging visual perception capabilities of Vision-Language Models (VLMs) to provide judgments that are close to human-level performance Lee et al. (2024a); Lu et al. (2024). To develop a framework capable of designing, evaluating, and self-improving in a cost-effective, rapid, and automated manner, we pose the following research question:
@@
In this work, we propose See2Refine that enables LLM designers to autonomously design, evaluate, and iteratively refine eHMI actions, utilizing visual perception from VLMs as automated feedback in a cost-effective manner. We position this work as an early-stage, controlled-scenario study of simulation-based eHMI refinement. In our experiments, we trained separate DesignerLLM models for each of the three eHMI modalities: lightbar, eyes, and arm. After three rounds of iterative learning, the DesignerLLM models exhibited a clear alignment in preferences with VLM raters. Additionally, our sampling method to expand the action database effectively reduces annotation costs and time without compromising performance. Subsequently, we recruited 18 participants to evaluate the eHMI actions generated by five LLMs: two 7B models (one base model, one DesignerLLM) and three state-of-the-art models. Extensive results demonstrate that leveraging the visual perception capabilities of VLMs can also enhance human evaluation scores. The main contributions of this study can be summarized as follows:
  • We propose See2Refine, a closed-loop framework that integrates VLM-based perceptual evaluation as automated visual feedback to improve LLM-based eHMI action designers in controlled simulated settings.
  • We discuss an efficient database expansion strategy that combines importance-based scenario sampling and diverse action generation to improve refinement efficiency without sacrificing preference alignment.
  • We show through a human study that VLM-guided refinement can transfer to measurable improvements in human-perceived eHMI quality across the evaluated scenarios.
2 Related Work
@@
Existing eHMI action-planning mostly follows predefined, designer-authored rules per modality: text/icons use traffic-rules messages Eisele and Petzoldt (2022); Eisma et al. (2021); color/light use flashing patterns based on intuitive color associations Bazilinskyy et al. (2019); Dey et al. (2020b); and anthropomorphic cues (e.g., eyes, arm gestures) draw on nonverbal communication principles Mahadevan et al. (2018); Ochiai and Toyoshima (2011). Although effective, these manual designs are labor-intensive and hard to scale to complex real-world scenarios Gui et al. (2023); de Winter and Dodou (2022). Recent work uses LLMs for eHMI action planning via prompt engineering Xia et al. (2025), but this remains prompt-dependent and largely task-level, lacking mechanism-level understanding, requiring more advanced methods for higher-quality action planning.
@@
Reinforcement Learning from AI Feedback (RLAIF) replaces the collection of human preferences in RLHF by using stronger teacher models to generate reward signals Bai et al. (2022b); Lee et al. (2023). In multimodal settings, VLMs can be fine-tuned as rubric-following judges for images and videos, achieving human-level agreement and performance competitive with GPT-4V Lu et al. (2024); Chen et al. (2024). Systems such as Prometheus-Vision, WildVision-Arena, WildVision-Bench, and MLLM-as-a-Judge show that VLM judges can follow instructions and output pointwise or pairwise scores that align with human preference orderings Lee et al. (2024a); Lu et al. (2024); Chen et al. (2024), while VHELM and LLaVA-Critic further standardize multimodal evaluation and reduce reliance on closed models Lee et al. (2024b); Xiong et al. (2025). Multimodal variants of RLAIF and RLHF-V use AI-generated preferences to train reward models and optimize large VLM policies with minimal human labeling Ahn et al. (2024); Yu et al. (2025); Sun et al. (2023). For eHMI, VLM judges correlate well with human ratings on LLM-generated action clips, enabling closed-loop optimization where an LLM proposes actions, a VLM scores them, and updates (e.g., policy gradient or DPO-style) refine the space without human studies Wang et al. (2024).
@@ 3 Method
As shown in Figure 2, our See2Refine framework contains five components, including (1) Scenario-Message Pair Generation, (2) Model Asset & Action Rendering, (3) Multi-Metric Evaluation System, (4) Format-aware Fine-tuning, and (5) Iterative Preference-based Learning. In the following, we describe each component in detail.
@@ 3.1 Scenario-Message Pair Generation
Inspired by existing eHMI scenario studies Dey et al. (2020a); Colley and Rukzio (2020), we first generate traffic scenarios by combining six different factors. The factors we consider are listed below. By combining these factors, we generate a total of 6,912 unique condition combinations.
  • Communication Relationships (4): 1st-person and 3rd-person perspective $\times$ one-to-one and one-to-many.
@@
  • Message Type (8): Instruction, advisory, question, answer, current, historical, predictive, and warn.
@@
  • Safety Implications (3): Critical (0–5 m), moderate (5–10 m), and routine (>>10 m).
Second, conditioned on these scenarios, we generate 20,736 intended messages containing scenario information using larger LLMs, including OpenAI GPT-4.1 OpenAI (2025), Claude 3.7 Sonnet Anthropic (2025), and Qwen3-235B-A22B Qwen Team, Alibaba Cloud (2025). To eliminate redundancy, we employ a sentence transformer Reimers and Gurevych (2019) to compare messages. We sample 70% of these messages using the farthest point sampling algorithm Eldar et al. (1997) to ensure sufficient variance. In addition, we conduct a small user-rating study on 200 randomly selected scenario–message pairs. Two eHMI researchers independently rate the validity of each pair on a 7-point Likert scale (1=Strongly Disagree, 7=Strongly Agree). The mean rating is 5.3 (SD=1.9), between “Somewhat Agree” (5) and “Agree” (6), suggesting that the generated scenarios are generally of high quality.
@@ 3.2.1 Model Development and Definition
In this study, we focus on three eHMI modalities: lightbar, eyes, and arm. The first two modalities represent classic eHMI designs Dey et al. (2020a); Deb et al. (2018); Rouchitsas and Alm (2019); Benderius et al. (2017); Colley et al. (2020) with simpler control mechanisms, while the arm modality requires managing more parameters, which presents additional challenges for DesignerLLM. In Blender (Blender Foundation, 2025), we implement each modality in two message emitter types (self-driving cars and delivery robots), resulting in six 3D models, as shown in Figure 3 (a). Modality descriptions and permitted actions are as follows:
@@
  • Light Bar is mounted on the vehicle front, slightly extending to both wings. It has 16 binary lighting regions (0=off, 1=on); e.g., “0000000011111111” denotes left-half off and right-half on (vehicle perspective). The brightness is fixed and the color is cyan Dey et al. (2020b).
@@
We also specify transition speeds between action states. The eye and arm transitions range from 0.5 to 2.0s (0.1s steps), while the light bar transitions range from 0.1 to 1.0s (0.1s per step). This enables smooth sequences while preserving temporal detail in the video FPS.
@@ 3.2.2 Rendering Pipeline
For rendering, we typically consider three conditions: camera direction, camera distance (relative to the object), and background scenes, as illustrated in Figure 3 (b). Regarding the direction of the camera, aligned with Section 3.1, we create 12 cameras arranged clockwise. Regarding the camera distance, which reflects the safety implications (Section 3.1), we position the cameras at three different distances. These positions can be accessed via Python to generate videos that simulate the perspectives of other road users. We use a plain background, unlike the previous work Xia et al. (2025), to avoid contextual cues and reduce rendering time. Removing scene backgrounds decreases the rendering time to approximately $1/10$, while lower GPU memory usage also enables more parallel threads, further reducing wall-clock time by around 4× for batch rendering.
We render action videos using Blender version 4.5 Blender Foundation (2025) on a GPU-equipped device featuring four NVIDIA GTX Titan cards with 24 GB of RAM each, allowing up to 16 concurrent threads. The average rendering time for 1,000 clips is approximately one hour, with each thread consuming around 1.1 GB of GPU RAM. Each frame is rendered at a resolution of 512×512 pixels, centered on the AV equipped with eHMI, as shown in Figure 3 (c). These clips are rendered at four FPS, with a maximum action change rate for each eHMI set to ensure that the frame rate captures all relevant details, following the Nyquist-Shannon sampling theorem Nyquist (1928).
@@ 3.3 Multi-Metric Evaluation System
Recent studies show that VLMs demonstrate capabilities and preferences similar to those of humans Lee et al. (2024a); Lu et al. (2024); Chen et al. (2024). In our task, our aim is to borrow knowledge from existing Human-Computer Interaction research Colley et al. (2020); Colley and Rukzio (2020); Colley et al. (2025) and use the features of VLMs to develop a multi-metric evaluation system.
@@ 3.3.1 Kernel Score: Composite Quality Metric
We use this kernel score as a simple default aggregation rather than as a claim of task-independent optimal weighting.
@@ 3.4 Format-aware Fine-tuning
This phase trains DesignerLLM-Apprentice to reliably output eHMI-specific action formats, as shown in Figure 2. We fine-tune two base model sizes, Qwen2.5-7B-Instruct and Qwen2.5-1.5B-Instruct Qwen Team (2024) as “apprentices” for later optimization. Training data is sampled from the shared action database by selecting the candidate action with the best kernel score per scenario.